# Two-Timescale Design for Movable Antennas Enabled-Multiuser MIMO Systems

### v1

@@ I Introduction
Despite the promising potential of MA technology, research in this field remains in its early stage, with most studies focusing on algorithmic design and theoretical performance limits. These studies generally assume that each antenna element can be adjusted globally and in real-time within the transmit/receive region, relying on the availability of instantaneous channel state information (CSI) [26]. However, such an approach presents several critical challenges. First, although mechanical MA systems, as discussed in [16] and [17], provide flexible movement in 2D and 3D spaces, this flexibility incurs substantial movement energy consumption due to the need for frequent adjustments to track optimal positions, compounded by delays resulting from mechanical response speed limitations. In contrast, fluid antenna systems (FAS), as studied in [12], [13], [27], are more compact and energy-efficient; however, their movement is confined to 1D spaces, limiting their adaptability in dynamic environments. Second, the reliance on complete and instantaneous CSI across the movement region to optimize MA positions poses an additional fundamental challenge. Most existing schemes rely on full CSI; however, acquiring accurate real-time CSI across the entire movement region is impractical and even infeasible, especially in fast-fading channels. This limitation hinders the practical deployment of MAs in real-world systems.
@@
    We further extend our analysis to zero-forcing (ZF) beamforming and propose an algorithm that resorts to AO, SCA, and majorization-minimization (MM) techniques. This iterative approach optimizes a surrogate function, which serves as a lower bound for the ergodic sum rate, to refine the antenna positions for enhanced performance.
@@
The rest of the paper is structured as follows: Section II introduces the system model and problem formulation. Section III presents the achievable rate derivation and antenna position optimization under MRT beamforming. Section IV derives the achievable rate and optimizes MAs’ positions under ZF beamforming. Section V presents the numerical results, and Section VI concludes the paper.
@@ II-A MA-enabled MU-MIMO System
where $\kappa_{m}\geq 0$ is the Rician factor, $\beta_{m}$ represents the large-scale fading coefficient, $\bar{\boldsymbol{h}}_{m}(\boldsymbol{t})\in\mathbb{C}^{N\times 1}$ denotes the deterministic LoS component, and $\tilde{\boldsymbol{h}}_{m}(\boldsymbol{t})\in\mathbb{C}^{N\times 1}$ represents the random non-line-of-sight (NLoS) component. Specifically, the entries of $\tilde{\boldsymbol{h}}_{m}$ are independent and identically distributed (i.i.d) circularly symmetric complex Gaussian random variables with zero mean and unit variance. To describe the LoS component, let $\theta_{m}\in\left[-\frac{\pi}{2},\frac{\pi}{2}\right]$ and $\phi_{m}\in\left[-\frac{\pi}{2},\frac{\pi}{2}\right]$ represent the elevation and azimuth AoDs for user $m$, respectively. The propagation distance difference of the LoS path between the MA position $\boldsymbol{t}$ and the reference position $\boldsymbol{o}$ is given by $\boldsymbol{t}_{n}^{T}\left[\cos\theta_{m}\sin\phi_{m},\sin\theta_{m}\right]^{T}\triangleq\boldsymbol{t}_{n}^{T}\boldsymbol{a}_{m}$, and the phase difference of the LoS path is $\frac{2\pi}{\lambda}\boldsymbol{t}_{n}^{T}\boldsymbol{a}_{m}$, with $\lambda$ denoting the carrier wavelength. Thus, the channel vector of the LoS component between the BS and $m$-th user can be expressed as
@@ II-B Two-Timescale Transmission Scheme
Building on this insight, we propose a hierarchical two-timescale transmission scheme, as shown in Fig. 2. In the first stage, the antenna positions at the BS are optimized based on statistical CSI, leveraging the stability of the LoS component to enhance long-term system performance. Once the optimal antenna positions are determined and fixed, the system transitions to the small timescale phase, where instantaneous CSI is acquired using conventional channel estimation techniques in multiple-antenna systems. This instantaneous CSI is then used to design the downlink beamforming, ensuring that the dynamic NLoS channel fluctuations are addressed in real time.
Compared to schemes requiring continuous MA repositioning based on instantaneous CSI, our two-timescale design effectively reduces the update frequency of MAs’ positions, lowering channel estimation overhead and movement energy consumption. This is particularly beneficial in environments with rapid NLoS fluctuations, where real-time MA adjustments are impractical. By integrating the NLoS component into the beamforming process after antenna positions are fixed, our method preserves the performance gains from this random component, combining the advantages of both statistical and instantaneous CSI. This two-timescale approach thus strikes an effective balance between practicality and performance.
@@ II-C Problem Formulation
Under the proposed two-timescale transmission framework, the objective is to maximize the ergodic sum rate for all users by jointly optimizing the antenna positions $\{{\boldsymbol{t}_{n}}\}$ in the large timescale and transmit beamforming vectors $\{{\boldsymbol{w}_{m}}\}$ in the small timescale at the BS. This leads to the formulation of a general optimization problem as follows
@@
[…]
@@
where ${\mathbb{E}}\left[\cdot\right]$ in (6a) represents the expectation over all possible channel realizations, $P_{\text{tot}}\geq 0$ in (6b) denotes the maximum transmit power of the BS, (6c) ensures a minimum separation distance between any two adjacent MAs to avoid coupling effects, and (6d) specifies the feasible region for MAs’ positions.
@@ III-A Ergodic Rate Analysis with MRT beamforming
$\quad\quad\quad\quad\quad R_{\text{MRT},m}^{app}=\log_{2}\left(1+\frac{\beta_{m}^{2}\left(\frac{2N\kappa_{m}+N}{\left(\kappa_{m}+1\right)^{2}}+N^{2}\right)}{\sum_{j=1,j\neq m}^{M}{\beta_{m}\beta_{j}\frac{\kappa_{m}\kappa_{j}\left|\boldsymbol{\bar{h}}_{j}\left(\boldsymbol{t}\right)\boldsymbol{\bar{h}}_{m}^{H}\left(\boldsymbol{t}\right)\right|^{2}+N\left(\kappa_{m}+\kappa_{j}+1\right)}{\left(\kappa_{m}+1\right)\left(\kappa_{j}+1\right)}}+\frac{\sigma_{m}^{2}}{P}N\sum_{m=1}^{M}{\beta_{m}}}\right).$ (13)
$\overline{\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ }$
Remark 1: From (13), we can draw three key insights. First, under MRT beamforming in the two-timescale design, MA does not yield any ergodic rate gain through antenna position tuning when $M=1$, as such gain stems from interference mitigation or channel decorrelation between $\boldsymbol{\bar{h}}_{j}\left(\boldsymbol{t}\right)$ and $\boldsymbol{\bar{h}}_{m}^{H}\left(\boldsymbol{t}\right)$, $\forall j\neq m$. Second, with $\kappa_{m}=0$ in the Rician fading channel, the absence of a deterministic LoS component will render MA ineffective in reshaping the channel since the summation term in the denominator of (13) vanishes, and thus $R_{\text{MRT},m}^{app}$ no longer related to $\boldsymbol{t}$. Third, as the number of antennas $N$ becomes asymptotically large, i.e., $N\rightarrow\infty$, the performance gain of MA diminishes due to the channel hardening, where FPAs already achieve asymptotically orthogonal channels for different users, i.e., $\big{|}\frac{1}{N}\boldsymbol{\bar{h}}_{j}\left(\boldsymbol{t}_{\text{FPA}}\right)\boldsymbol{\bar{h}}_{m}^{H}\left(\boldsymbol{t}_{\text{FPA}}\right)\big{|}^{2}\rightarrow 0,\forall j,m$.
@@ III-C Constructing a Surrogate Function via SCA Technique
×$\qquad\qquad\quad\times\left(b_{m}\left(\boldsymbol{t}_{n}\right)-b_{m}\left(\boldsymbol{t}_{n}^{\ell}\right)\right),$ (19)
@@
×$\times(\sin\theta_{m}-\sin\theta_{j})\big{|},$ (26b)
@@
With the above transformation, we construct a convex upper-bound surrogate function for $b_{m}(\boldsymbol{t}_{n}^{\ell})$ and, consequently, derive a concave lower-bound surrogate function for (18). This approach, combined with the first-order approximation in (19), facilitates the design of an efficient antenna position optimization algorithm that iteratively refines the solution within the SCA framework. The sole remaining hurdle in solving the problem (P2.n) lies in the non-convex minimum distance constraint (6d), which is equivalent to ensuring that $\geq D_{\min}^{2},\forall n,i\in\mathcal{N},i\neq n$. Given that $\lVert\boldsymbol{t}_{n}-\boldsymbol{t}_{i}\rVert^{2}$ is convex with respect to $\boldsymbol{t}_{n}$, it is lower bounded by its first-order Taylor expansion as follows
@@ III-D Overall Algorithm and Computational Complexity Analysis
×$\!\!\times\big{(}\nabla b_{m}\left(\boldsymbol{t}_{n}^{\ell}\right)^{T}\left(\boldsymbol{t}_{n}-\boldsymbol{t}_{n}^{\ell}\right)+\frac{\psi_{m,n}}{2}\lVert\boldsymbol{t}_{n}-\boldsymbol{t}_{n}^{\ell}\rVert^{2}\big{)}\!\!\!$ (31a)
[…]
@@
The overall AO-based algorithm is detailed in Algorithm 1, deriving a solution for the problem (P2) by iteratively solving (P3.n). The algorithm guarantees convergence to local optimal points since the objective value is non-decreasing in each iteration. Given that the problem is a QCP, the overall computational complexity of Algorithm 1 is $\mathcal{O}(L_{1}\ln\frac{1}{\varepsilon}N^{4.5})$ [37], where $L_{1}$ represents the number of iterations required for convergence, and $\varepsilon$ denotes the prescribed accuracy.
@@
4:     For each iteration, solve the problem (P3.n) for $N$ antennas using convex optimization methods obtain updated antenna positions $\boldsymbol{t}_{n}^{\ell+1}$.

### v2

# Two-Timescale Design for Movable Antenna-Enabled Multiuser MIMO Systems

@@ I Introduction
Despite the promising potential of MA technology, research in this field remains in its early stage, with most studies focusing on algorithmic design and theoretical performance limits. These studies generally assume that each antenna element can be adjusted globally and in real-time within the transmit/receive region, relying on the availability of instantaneous channel state information (CSI) [26]. However, such an approach presents several critical challenges. First, although mechanical MA systems, as discussed in [16] and [17], provide flexible movement in 2D and 3D spaces, this flexibility incurs substantial movement energy consumption due to the need for frequent adjustments to track optimal positions, compounded by delays resulting from mechanical response speed limitations. Fluid antenna systems (FAS) can be more compact and energy-efficient [12], [13], [27]. Second, the reliance on complete and instantaneous CSI across the movement region to optimize MA positions poses an additional fundamental challenge. Most existing schemes rely on full CSI; however, acquiring accurate real-time CSI across the entire movement region is impractical and even infeasible, especially in fast-fading channels. This limitation hinders the practical deployment of MAs in real-world systems.
@@
    We further analyze zero-forcing (ZF) beamforming and propose an algorithm that resorts to AO, SCA, and majorization-minimization (MM) techniques. This iterative approach optimizes a surrogate function, which serves as a lower bound for the ergodic sum rate, to refine the antenna positions for enhanced performance. In addition, we incorporate the effect of spatial correlation into the channel model and analyze its impact on the ergodic rate for both cases.
@@
It is worth emphasizing that the proposed two-timescale design offers distinct advantages across different deployment scenarios. In semi-static environments such as industrial IoT, smart home, and fixed wireless access, the framework achieves substantial and stable performance gains with dramatically reduced CSI overhead and minimal MA updates, making it highly practical and effective. For mobile users, although frequent statistical CSI and MA position updates may be required to maintain performance, the design still effectively reduces instantaneous CSI overhead and provides a flexible fallback to FPA-based configurations when positioning inaccuracies become significant. This adaptability ensures the practicality of the proposed scheme under a wide range of network conditions.
The rest of the paper is structured as follows: Section II introduces the system model and problem formulation. Section III presents the achievable rate derivation and antenna position optimization under MRT beamforming. Section IV derives the achievable rate and optimizes MAs’ positions under ZF beamforming. Section V derives the achievable rate with spatial correlation. Section VI presents the numerical results, and Section VII concludes the paper.
@@ II-A MA-enabled MU-MIMO System
where $\kappa_{m}\geq 0$ is the Rician factor, $\beta_{m}$ represents the large-scale fading coefficient, $\bar{\boldsymbol{h}}_{m}(\boldsymbol{t})\in\mathbb{C}^{N\times 1}$ denotes the deterministic LoS component, and $\tilde{\boldsymbol{h}}_{m}(\boldsymbol{t})\in\mathbb{C}^{N\times 1}$ represents the random non-line-of-sight (NLoS) component. Specifically, the NLoS component $\tilde{\boldsymbol{h}}_{m}(\boldsymbol{t})$ should be modeled as a spatially correlated Gaussian process with a covariance matrix $\tilde{\boldsymbol{h}}_{m}(\boldsymbol{t})=\boldsymbol{S}(\boldsymbol{t})\boldsymbol{u}_{m}$ with the Gaussian random variables $\boldsymbol{u}_{m}\sim\mathcal{C}\mathcal{N}\left(0,\boldsymbol{I}\right)$, and the spatial correlation matrix $\boldsymbol{S}\left(\boldsymbol{t}\right)$, which captures spatial dependencies due to the scattering environment. A widely adopted model for $\boldsymbol{S}(\boldsymbol{t}$ is the Bessel function model [12],[13], where the correlation between two antennas is expressed as $\left[\boldsymbol{S}(\boldsymbol{t})\right]_{n,n^{\prime}}=J_{0}\left(\frac{2\pi\lVert\boldsymbol{t}_{n}-\boldsymbol{t}_{n^{\prime}}\rVert}{\lambda}\right),\forall n,n^{\prime}\in\mathcal{N}$, where $J_{0}(\cdot)$ is the zero-order Bessel function of the first kind, $\lVert\boldsymbol{t}_{n}-\boldsymbol{t}_{n^{\prime}}\rVert$ is the distance between $n$-th and $n^{\prime}$-th elements on the antenna array. For analytical tractability in our study, we simplify the NLoS component to i.i.d. complex Gaussian variables with zero mean and unit variance, approximating $\boldsymbol{S}(\boldsymbol{t})$ as an identity matrix. This simplification is supported by the imposed minimum antenna spacing constraint, which reduces spatial correlation, with a reasonable approximation error. We extend this analysis in Sections V and VI to evaluate the impact of spatial correlation, and in the following analysis and design, we simplify the NLoS modeling by assuming that the entries of $\tilde{\boldsymbol{h}}_{m}$ are independent and identically distributed (i.i.d.) circularly symmetric complex Gaussian random variables with zero mean and unit variance. To describe the LoS component, let $\theta_{m}\in\left[-\frac{\pi}{2},\frac{\pi}{2}\right]$ and $\phi_{m}\in\left[-\frac{\pi}{2},\frac{\pi}{2}\right]$ represent the elevation and azimuth AoDs for user $m$, respectively. The propagation distance difference of the LoS path between the MA position $\boldsymbol{t}$ and the reference position $\boldsymbol{o}$ is given by $\boldsymbol{t}_{n}^{T}\left[\cos\theta_{m}\sin\phi_{m},\sin\theta_{m}\right]^{T}\triangleq\boldsymbol{t}_{n}^{T}\boldsymbol{a}_{m}$, and the phase difference of the LoS path is $\frac{2\pi}{\lambda}\boldsymbol{t}_{n}^{T}\boldsymbol{a}_{m}$, with $\lambda$ denoting the carrier wavelength. Thus, the channel vector of the LoS component between the BS and $m$-th user can be expressed as
@@ II-B Two-Timescale Transmission Scheme
Building on this insight, we propose a hierarchical two-timescale transmission scheme, as shown in Fig. 2. In the first stage, the antenna positions at the BS are optimized based on statistical CSI, leveraging the stability of the LoS component to enhance long-term system performance [^1] Incorporating ray tracing–based predictive CSI earlier, more precise environmental data refine position-based CSI, reduces CSI estimation overhead, and strengthens our two-timescale MA framework without altering its design.. Once the optimal antenna positions are determined and fixed, the system transitions to the small timescale phase, where instantaneous CSI is acquired using conventional channel estimation techniques in multiple-antenna systems. This instantaneous CSI is then used to design the downlink beamforming, ensuring that the dynamic NLoS channel fluctuations are addressed in real time.
Compared to schemes requiring continuous MA repositioning based on instantaneous CSI, our two-timescale design effectively reduces the update frequency of MAs’ positions [^2] This design is particularly beneficial in environments with rapid NLoS fluctuations but stable LoS information, while in high-mobility scenarios, the MA positions may still need to be updated more frequently based on changing statistical CSI to maintain performance. However, the two-timescale design continues to offer practical advantages by significantly reducing instantaneous CSI overhead.., lowering channel estimation overhead and movement energy consumption [^3] For a quantitative energy trade-off analysis, we model each MA’s repositioning energy as $E_{n}=P_{x}\frac{|\Delta x_{n}|}{v_{x}}+P_{y}\frac{|\Delta y_{n}|}{v_{y}}$ [43], where $P_{x}=P_{y}=8\,\mathrm{W}$ (typical stepper motor power [44]) and $v_{x}=v_{y}\approx 0.94\,\mathrm{mm/ms}$ (based on a 10 mm lead screw [43]). In an MA array of $N=6$ with an average movement of $|\Delta x_{n}|=|\Delta y_{n}|=\lambda/2=25\,\mathrm{mm}$ at 12 GHz, repositioning consumes $E_{\mathrm{MA}}\approx 1.28\,\mathrm{J}$ per cycle, whereas data transmission at $P_{\mathrm{TR}}=0.1\,\mathrm{W}$ over $300\,\mathrm{ms}$ uses only $E_{\mathrm{TR}}=0.03\,\mathrm{J}$, making $E_{\mathrm{MA}}$ 42 times higher if executed every frame. By reducing the repositioning frequency (e.g., from every 300 ms to every 3 s) under stable statistical CSI, a two-timescale framework cuts $E_{\mathrm{MA}}$ by a factor of 10 while preserving performance.. By integrating the NLoS component into the beamforming process after antenna positions are fixed, our method preserves the performance gains from this random component, combining the advantages of both statistical and instantaneous CSI. This two-timescale approach thus strikes an effective balance between practicality and performance.
@@ II-C Problem Formulation
Under the proposed two-timescale transmission framework, the objective is to maximize the ergodic sum rate for all users by jointly optimizing the antenna positions $\{{\boldsymbol{t}_{n}}\}$ in the large timescale and transmit beamforming vectors $\{{\boldsymbol{w}_{m}}\}$ in the small timescale at the BS. This leads to the formulation of a general optimization problem as follows [^4] The two-timescale framework can be readily extended to multi-cell networks, where inter-cell interference becomes a key consideration, and remains sufficiently general to be adapted for various advanced wireless scenarios, e.g., machine-type communication, wireless sensing, and indoor localization.
@@
$\text{s}.\text{t}.$ $\,\,\sum_{m=1}^{M}{\lVert\boldsymbol{w}_{m}\rVert^{2}\leq P_{\text{tot}}},\,\,$ (6b)
@@
where $\mathbb{E}$$\left$$\cdot\right$] in (6a) represents the expectation over all possible channel realizations, $P_{\text{tot}}\geq 0$ in (6b) denotes the maximum transmit power of the BS, (6c) ensures a minimum separation distance between any two adjacent MAs to avoid coupling effects and mitigate spatial correlation, and (6d) specifies the feasible region for MAs’ positions.
@@ III-A Ergodic Rate Analysis with MRT beamforming
$\quad\quad\quad\quad\quad R_{\text{MRT},m}^{app}=\log_{2}\Bigg(1+\frac{\beta_{m}^{2}\left(\frac{2N\kappa_{m}+N}{\left(\kappa_{m}+1\right)^{2}}+N^{2}\right)}{\sum_{j=1,j\neq m}^{M}{\beta_{m}\beta_{j}\frac{\kappa_{m}\kappa_{j}\left|\boldsymbol{\bar{h}}_{j}\left(\boldsymbol{t}\right)\boldsymbol{\bar{h}}_{m}^{H}\left(\boldsymbol{t}\right)\right|^{2}+N\left(\kappa_{m}+\kappa_{j}+1\right)}{\left(\kappa_{m}+1\right)\left(\kappa_{j}+1\right)}}+\frac{\sigma_{m}^{2}}{P}N\sum_{m=1}^{M}{\beta_{m}}}\Bigg).$ (13)
Remark 1: From (13), we can draw three key insights. First, under MRT beamforming in the two-timescale design, when $M=1$ there is no ergodic rate gain from antenna position tuning. This is because the gain originates from interference mitigation or channel decorrelation between $\boldsymbol{\bar{h}}_{j}\left(\boldsymbol{t}\right)$ and $\boldsymbol{\bar{h}}_{m}^{H}\left(\boldsymbol{t}\right)$, $\forall j\neq m$, which is achieved by optimizing $\boldsymbol{t}$ to tune the phases of the LoS components and reduce $\left|\boldsymbol{\bar{h}}_{j}\left(\boldsymbol{t}\right)\boldsymbol{\bar{h}}_{m}^{H}\left(\boldsymbol{t}\right)\right|^{2}$. Second, with $\kappa_{m}=0$ in the Rician fading channel, the absence of a deterministic LoS component will render MA ineffective in reshaping the channel since the summation term in the denominator of (13) vanishes, and thus $R_{\text{MRT},m}^{app}$ no longer related to $\boldsymbol{t}$. Third, as the number of antennas $N$ becomes asymptotically large, i.e., $N\rightarrow\infty$, the performance gain of MA diminishes due to the channel hardening, where FPAs already achieve asymptotically orthogonal channels for different users, i.e., $\big|\frac{1}{N}\boldsymbol{\bar{h}}_{j}\left(\boldsymbol{t}_{\text{FPA}}\right)\boldsymbol{\bar{h}}_{m}^{H}\left(\boldsymbol{t}_{\text{FPA}}\right)\big|^{2}\rightarrow 0,\forall j,m$.
@@ III-C Constructing a Surrogate Function via SCA Technique
$\qquad\qquad\quad\times\left(b_{m}\left(\boldsymbol{t}_{n}\right)-b_{m}\left(\boldsymbol{t}_{n}^{\ell}\right)\right),$ (19)
@@
$\times(\sin\theta_{m}-\sin\theta_{j})\big|,$ (26b)
@@
With the above transformation, we construct a convex upper-bound surrogate function for $b_{m}(\boldsymbol{t}_{n}^{\ell})$ and, consequently, derive a concave lower-bound surrogate function for (18). This approach, combined with the first-order approximation in (19), facilitates the design of an efficient antenna position optimization algorithm that iteratively refines the solution within the SCA framework. The sole remaining hurdle in solving the problem (P2.n) lies in the non-convex minimum distance constraint (6d), which is equivalent to ensuring that $\lVert\boldsymbol{t}_{n}-\boldsymbol{t}_{i}\rVert^{2}\geq D_{\min}^{2},\forall n,i\in\mathcal{N},i\neq n$. Given that $\lVert\boldsymbol{t}_{n}-\boldsymbol{t}_{i}\rVert^{2}$ is convex with respect to $\boldsymbol{t}_{n}$, it is lower bounded by its first-order Taylor expansion as follows
@@ III-D Overall Algorithm and Computational Complexity Analysis
$\!\!\times\big(\nabla b_{m}\left(\boldsymbol{t}_{n}^{\ell}\right)^{T}\left(\boldsymbol{t}_{n}-\boldsymbol{t}_{n}^{\ell}\right)+\frac{\psi_{m,n}}{2}\lVert\boldsymbol{t}_{n}-\boldsymbol{t}_{n}^{\ell}\rVert^{2}\big)\!\!\!$ (31a)
$\,\text{s.t.}\,\,\mathcal{T}^{\text{lb},\ell}(\boldsymbol{t}_{n})\geq D_{\min}^{2},\forall n,i\in\mathcal{N},$ (31b)
@@
The overall AO-based algorithm is detailed in Algorithm 1, deriving a solution for the problem (P2) by iteratively solving (P3.n). The algorithm guarantees convergence to local optimal points since the objective value is non-decreasing in each iteration. The total computational complexity is $\mathcal{O}(L_{1}(N^{2.5}\log\frac{1}{\varepsilon}+NM^{2}+MN^{2}))$, derived as follows: for each antenna, a convex quadratic constrained program (QCP) is solved using an interior-point method, with complexity $\mathcal{O}(N^{1.5}\log\frac{1}{\varepsilon})$, where $\varepsilon$ denotes the prescribed accuracy. This accounts for the 2D optimization variable $\boldsymbol{t}_{n}=\left[x_{n},y_{n}\right]^{T}$ with $\mathcal{O}(N)$ convexified distance constraints and linear region constraints. Per iteration, evaluate the surrogate objective costs $\mathcal{O}(M^{2}+MN)$, calculate the gradient costs $\mathcal{O}(M^{2})$, and determine the Hessian bound costs $\mathcal{O}(M^{2})$. The algorithm iterates over all $N$ antennas $L_{1}$ times until convergence, summing to $\mathcal{O}(L_{1}(N^{2.5}\log\frac{1}{\varepsilon}+NM^{2}+MN^{2}))$.
@@
4:  For each iteration, solve the problem (P3.n) for NN antennas using convex optimization methods obtain updated antenna positions $\boldsymbol{t}_{n}^{\ell+1}$.